# Symplectic self-orthogonal quasi-cyclic codes

### v4

@@ II-A Linear codes
Let $\mathbb{F}_{q}$ be a finite field with $q$ elements, where $q$ is a prime power. Let $\mathbb{F}_{q}^{*}=\mathbb{F}_{q}\setminus\{0\}$. An $\left[\ell n,k\right]_{q}$ linear code $\mathscr{C}$ is a $k$-dimensional subspace of $\mathbb{F}_{q}^{\ell n}$. For a codeword $\mathbf{c}=(c_{0},c_{1},\ldots,c_{\ell n-1})$ of $\mathscr{C}$, the Hamming weight of $\mathbf{c}$ is $\mathbf{w}_{H}(\mathbf{c})=|\left\{i\mid c_{i}\neq 0,0\leq i\leq\ell n-1\right\}|$. If $\ell n$ is even and $N=\ell n/2$, then the symplectic weight of $\mathbf{c}$ is $\mathbf{w}_{s}(\mathbf{c})=|\{i\mid(u_{i},u_{N+i})\neq(0,0),0\leq i\leq N-1\}|$. The support of $\mathbf{c}$, denoted by $\mathrm{Supp}(\mathbf{c})$, is the set of coordinates of nonzero entries in $\mathbf{c}$, which is a subset of $\{1,\ldots,\ell n\}$. If the minimum Hamming (resp. symplectic) weight of $\mathscr{C}$ is $d_{H}=\min\left\{\mathbf{w}_{H}(\mathbf{u})\mid\mathbf{u}\in\mathscr{C},\mathbf{u}\neq\mathbf{0}\right\}$ (resp. $d_{s}=\min\left\{\mathbf{w}_{s}(\mathbf{u})\mid\mathbf{u}\in\mathscr{C},\mathbf{u}\neq\mathbf{0}\right\}$), then $\mathscr{C}$ can be written as an $\left[\ell n,k,d_{H}\right]_{q}$ (resp. $\left[\ell n,k,d_{s}\right]_{q}^{s}$ ) code.
@@ II-B Cyclic codes and quasi-cyclic codes
Then, we have $\pi(xc(x))=\tau_{1}(\mathbf{c})$. Therefore, when studying cyclic codes, $\left[c(x)\right]$ is identified with the coefficient vector $\mathbf{c}$. Since a cyclic code is an ideal in the quotient ring $\mathbb{R}_{q,n}$ and $\mathbb{R}_{q,n}$ is principal, every cyclic code can be generated by a single polynomial in $\mathbb{R}_{q,n}$. For an $\left[n,k\right]_{q}$ cyclic code $\mathscr{C}$, we call the monic polynomial $g(x)$ with least degree $n-k$ that generates $\mathscr{C}$ as the generator polynomial of $\mathscr{C}$. We use $d_{H}(\mathscr{C})$ or $d(\left[g(x)\right])$ to denote the minimum Hamming distance of $\mathscr{C}$. The parity check polynomial of $\mathscr{C}$ is $h(x)=(x^{n}-1)/g(x)$. The Euclidean dual of $\mathscr{C}$ is also a cyclic code, which has generator polynomial $g^{\perp_{e}}(x)=h_{0}^{-1}x^{\deg(h(x))}h(x^{-1})$, where $h_{0}$ is the constant term coefficient of $h(x)$. If $g^{\perp_{e}}(x)\mid g(x)$, then $\mathscr{C}$ is a Euclidean self-orthogonal code.
@@ II-C Quantum error-correction codes
Note: Throughout the paper, we agree on the following notation. All the indices $\ell$ of QC codes in the subsequent texts will be even and set $m=\ell/2$. For $f(x)=f_{0}+f_{1}x+\cdots+f_{n-1}x^{n-1}\in\mathbb{R}_{q,n}$, denote $\bar{f}(x)=\sum\limits_{i=0}^{n-1}{f_{i}x^{n-i}}$. $\mathbf{0}$ and $\mathbf{0}_{n}$ denote a appropriate zero matrix and a zero vector of length $n$, respectively. Additionally, we denote $\Omega_{mn}=\left(\begin{array}[]{cc}\mathbf{0}&I_{mn}\\ -I_{mn}&\mathbf{0}\end{array}\right)$, where $I_{mn}$ is the identity matrix of order $mn$.
@@ III The necessary and sufficient conditions for QC codes with index even to be symplectic self-orthogonal
Let $g(x)$ be a polynomial in $\mathbb{R}_{q,n}$ that divides ${{x}^{n}}-1$, and let $f_{j}(x)$ be polynomials in $\mathbb{R}_{q,n}$ for $11\leq j\leq\ell-$ 1, such that the greatest common divisor of $f_{0}(x),f_{1}(x),\ldots,f_{\ell-1}(x)$ and $\frac{x^{n}-1}{g(x)}$ is equal to $1$. If $\mathscr{C}$ is a QC code generated by $(\left[g(x){f}_{0}(x)\right]$, $\left[g(x){f}_{1}(x)\right],\ldots,\left[g(x){f}_{\ell-1}(x)\right])\in\mathbb{R}_{q,n}^{\ell}$, then $\mathscr{C}$ is called a $1$-generator QC code with index $\ell$. The generator matrix $G$ of $\mathscr{C}$ has the following form:
@@
Let $g_{i}(x)\mid({{x}^{n}}-1)$ and $f_{i,j}(x)$ be polynomials in $\mathbb{R}_{q,n}$ such that the greatest common divisor of $f_{i,0}(x),$ $f_{i,1}(x),\ldots,f_{i,{\ell-1}}(x)$ and $\frac{x^{n}-1}{g_{i}(x)}$ is equal to $1$, where $1\leq i\leq h$ and $10\leq j\leq\ell-$ 1. If $\mathscr{C}$ is a QC code with $h$ generators given by: $(\left[g_{1}(x)f_{1,0}(x)\right],\left[g_{1}(x)f_{1,1}(x)\right],\ldots,\left[g_{1}(x)f_{1,\ell-1}(x)\right])$, $(\left[g_{2}(x)f_{2,0}(x)\right],$ $\left[g_{2}(x)f_{2,1}(x)\right],\ldots,\left[g_{2}(x)f_{2,\ell-1}(x)\right]),\ldots$, $(\left[g_{h}(x)f_{h,0}(x)\right],$ $\left[g_{h}(x)f_{h,1}(x)\right],\ldots,\left[g_{h}(x)f_{h,\ell-1}(x)\right])\in\mathbb{R}_{q,n}^{\ell}$, then $\mathscr{C}$ is an $h$-generator QC code with index $\ell$, whose generator matrix is
@@
If $\mathbf{c}^{\prime}=\mathbf{0}_{n}$ or $\mathbf{c}^{\prime\prime}=\mathbf{0}_{n}$, then we have $a(x)g(x)f_{0}(x)\equiv 0\pmod{x^{n}-1}$ or $a(x)g(x)f_{1}(x)\equiv 0\pmod{x^{n}-1}$. By Lemma 5, $d_{s}(\mathbf{c})=d_{H}(\mathscr{C}_{3})$ or $d_{H}(\mathscr{C}_{4})$.
@@
Secondly, if there is $\alpha$ in $\mathbb{F}_{q}^{*}$ such that $\mathcal{I}^{(\alpha)}(\mathscr{C})=\gcd(f_{0}(x)+\alpha f_{1}(x),h(x))\neq 1$, i.e., $\mathcal{S}\neq\emptyset$, then $\mathrm{Supp}(\mathbf{c}^{\prime})=\mathrm{Supp}(\mathbf{c}^{\prime\prime})$ may occur. For $q=2$, if $\mathrm{Supp}(\mathbf{c}^{\prime})=\mathrm{Supp}(\mathbf{c}^{\prime\prime})$, then we have $\mathbf{c}^{\prime}=\mathbf{c}^{\prime\prime}$ $\in\mathscr{C}_{1}\cap\mathscr{C}_{2}$. Therefore, $d_{s}(\mathscr{C})\geq d_{H}(\left[g(x)\operatorname{lcm}(f_{0}(x),f_{1}(x))\right])$. For $q>2$, we can only determine $d_{s}(\mathscr{C})\geq\max\{d_{H}(\mathscr{C}_{1}),d_{H}(\mathscr{C}_{2})\}$.
@@
With the above notation, let $\mathscr{C}$ be a $1$-generator QC code with generator $(\left[f_{0}(x)g(x)\right],\left[f_{1}(x)g(x)\right])$. If $\gcd(\bar{f}_{0}(x),g(x))=1$, then the symplectic dual code $\mathscr{C}^{\perp_{s}}$ of $\mathscr{C}$ is a 2-generator QC code, whose generator matrix is
@@
where $(F_{0},F_{1})$ and $(\mathbf{0},G_{g^{\perp_{e}}})$ are the generator matrices of $1$-generator QC codes determined by $(\left[\bar{f}_{0}(x)\right],\left[\bar{f}_{1}(x)\right])$ and $(\left[0\right],\left[G_{g^{\perp_{e}}}\right])$, respectively.
@@
Let $\mathscr{C}^{\circ}$ denote the code generated by $G^{\perp_{s}}$. Then any codeword of $\mathscr{C}$ and $\mathscr{C}^{\circ}$ has the form $\mathbf{c_{1}}=(\left[a(x)f_{0}(x)g(x)\right],$ $\left[a(x)f_{1}(x)g(x)\right])$ and $\mathbf{c_{2}}=(\left[b(x)\bar{f}_{0}(x)\right],\left[b(x)\bar{f}_{1}(x)+c(x)g^{\perp_{e}}(x)\right])$, where $a(x),b(x),c(x)\in\mathbb{R}_{q,n}$. Using a similar approach of Theorem 2, it is easily proven that $\left\langle\mathbf{c_{1}},\mathbf{c_{2}}\right\rangle_{s}=\mathbf{c_{1}}\cdot\Omega_{mn}\cdot\mathbf{c_{2}}^{T}=0$. Therefore, $\mathscr{C}$ and $\mathscr{C}^{\circ}$ are mutually symplectic orthogonal.
@@
According to Definition 1, $\gcd(\bar{f}_{0}(x),\bar{f}_{1}(x),g^{\perp_{e}}(x))=1$. Given that $\gcd(\bar{f}_{0}(x),g(x))=1$, we have $\gcd(\bar{f}_{0}(x),\bar{f}_{1}(x),x^{n}-1)=1$ and $\operatorname{lcm}\left(g^{\perp_{e}}(x),\frac{x^{n}-1}{\gcd(x^{n}-1,\bar{f}_{0}(x))}\right)\equiv 0\pmod{x^{n}-1}$, i.e., the rank of $(F_{0},F_{1})$ is $n$ and the rows of $B$ and $G_{g^{\perp_{e}}}$ are linearly independent. Therefore, the rank of $G_{1}^{\perp_{s}}$ is $n+\deg(g(x))$. Thus, $\mathscr{C}$ and $\mathscr{C}^{\circ}$ are symplectic dual codes. ∎
@@
With the above notation, let $\mathscr{C}$ be a $1$-generator QC code with generator $(\left[f_{0}(x)g(x)\right],\left[f_{1}(x)g(x)\right])$ satisfying $\gcd(\bar{f}_{0}(x),g(x))=1$. For $\alpha\in\mathbb{F}_{q}^{*}$, define $\mathcal{I}^{(\alpha)}_{\perp}(\mathscr{C})=\gcd(\bar{f}_{1}(x)+\alpha\bar{f}_{0}(x),g^{\perp_{e}}(x))$. Let $\mathcal{S}_{\perp}=\{\mathcal{I}^{(\alpha)}_{\perp}(\mathscr{C})\mid\mathcal{I}^{(\alpha)}_{\perp}(\mathscr{C})\neq 1,\forall\alpha\in\mathbb{F}_{q}^{*}\}$. Suppose that $\mathscr{C}_{0}$, $\mathscr{C}_{1}$, $\mathscr{C}_{2}$, $\mathscr{C}_{3}$, $\mathscr{C}_{4}$, $\mathscr{C}_{5}$, $\mathscr{C}_{6}$ and $\mathscr{C}_{7}$ are cyclic codes determined by $\left[g^{\perp_{e}}(x)\right]$, $\left[\bar{f}_{0}(x)\right]$, $\left[\gcd(\bar{f}_{1}(x),g^{\perp_{e}}(x))\right]$, $\left[\frac{x^{n}-1}{\gcd(x^{n}-1,\bar{f}_{1}(x))}\right]$, $\left[\frac{x^{n}-1}{\gcd(x^{n}-1,\bar{f}_{0}(x))}\right]$, $\left[\gcd\left(\frac{x^{n}-1}{\gcd(x^{n}-1,\bar{f}_{0}(x))},g^{\perp_{e}}(x)\right)\right]$, $\left[\operatorname{lcm}(\bar{f}_{0}(x),\gcd(\bar{f}_{1}(x),g^{\perp_{e}}(x))\right]$ and $\left[\bar{f}_{0}(x)\frac{g^{\perp_{e}}(x)}{\gcd(\bar{f}_{1}(x),g^{\perp_{e}}(x))}\right]$, respectively.
@@
Let $t=(n-\deg(\gcd(x^{n}-1,\bar{f}_{1}(x)))-\deg(\gcd(x^{n}-1,\bar{f}_{0}(x))))$ and $\mathscr{C}_{8}$ be the cyclic code generated by $\left[\bar{f}_{1}(x)\right]$. Since $\gcd(\bar{f}_{0}(x),g(x))=1$, by Theorem 5, the symplectic dual of $\mathscr{C}$ is a 2-generator QC code with generator matrix
@@
If $a(x)g(x)f_{0}(x)\not\equiv 0\pmod{x^{n}-1}$ and $a(x)g(x)f_{1}(x)\not\equiv 0\pmod{x^{n}-1}$, there are the following scenarios to discuss.
@@
Thirdly, if there exists $\alpha$ in $\mathbb{F}_{q}^{*}$ such that $\gcd(\bar{f}_{1}(x)+\alpha\bar{f}_{0}(x),g^{\perp_{e}}(x))\neq 1$, i.e., $\mathcal{S}_{\perp}\neq\emptyset$, then there may exist $\mathrm{Supp}(\left[a(x)\bar{f}_{0}(x)\right])=\mathrm{Supp}(\left[a(x)\bar{f}_{0}(x)+b(x)g^{\perp_{e}}(x)\right])$. Thus, for $q>2$, we can only deduce $d_{s}(\mathscr{C}^{\perp_{s}})\geq\max\left\{d_{H}(\mathscr{C}_{1}),d_{H}(\mathscr{C}_{2})\right\}$. But for $q=2$, we have $d_{s}(\mathscr{C}^{\perp_{s}})\geq d_{H}(\left[\operatorname{lcm}(\bar{f}_{0}(x),\gcd(\bar{f}_{1}(x),g^{\perp_{e}}(x)))\right])$.
For the case of $\mathrm{Supp}(\left[a(x)\bar{f}_{0}(x)\right])\neq$ $\mathrm{Supp}(\left[a(x)\bar{f}_{0}(x)+\right.$ $\left.b(x)g^{\perp_{e}}(x)\right])$, we have Equation (25) holds.

### v5

@@ II-A Linear codes
Let $\mathbb{F}_{q}$ be a finite field with $q$ elements, where $q$ is a prime power. Let $\mathbb{F}_{q}^{*}=\mathbb{F}_{q}\setminus\{0\}$. An $\left[\ell n,k\right]_{q}$ linear code $\mathscr{C}$ is a $k$-dimensional subspace of $\mathbb{F}_{q}^{\ell n}$. For a codeword $\mathbf{c}=(c_{0},c_{1},\ldots,c_{\ell n-1})$ of $\mathscr{C}$, the Hamming weight of $\mathbf{c}$ is $\mathbf{w}_{H}(\mathbf{c})=|\left\{i\mid c_{i}\neq 0,0\leq i\leq\ell n-1\right\}|$. If $\ell n$ is even and $N=\ell n/2$, then the symplectic weight of $\mathbf{c}$ is $\mathbf{w}_{s}(\mathbf{c})=|\{i\mid(c_{i},c_{N+i})\neq(0,0),0\leq i\leq N-1\}|$. The support of $\mathbf{c}$, denoted by $\mathrm{Supp}(\mathbf{c})$, is the set of coordinates of nonzero entries in $\mathbf{c}$, which is a subset of $\{0,1,\ldots,\ell n-1\}$. If the minimum Hamming (resp. symplectic) weight of $\mathscr{C}$ is $d_{H}=\min\left\{\mathbf{w}_{H}(\mathbf{c})\mid\mathbf{c}\in\mathscr{C},\mathbf{c}\neq\mathbf{0}\right\}$ (resp. $d_{s}=\min\left\{\mathbf{w}_{s}(\mathbf{c})\mid\mathbf{c}\in\mathscr{C},\mathbf{c}\neq\mathbf{0}\right\}$), then $\mathscr{C}$ can be written as an $\left[\ell n,k,d_{H}\right]_{q}$ (resp. $\left[\ell n,k,d_{s}\right]_{q}^{s}$ ) code.
@@ II-B Cyclic codes and quasi-cyclic codes
Then, we have $\pi(xc(x))=\tau_{1}(\mathbf{c})$. Therefore, when studying cyclic codes, $\left[c(x)\right]$ is identified with the coefficient vector $\mathbf{c}$. Since a cyclic code is an ideal in the quotient ring $\mathbb{R}_{q,n}$ and $\mathbb{R}_{q,n}$ is principal, every cyclic code can be generated by a single polynomial in $\mathbb{R}_{q,n}$. For an $\left[n,k\right]_{q}$ cyclic code $\mathscr{C}$, we call the monic polynomial $g(x)$ with least degree $n-k$ that generates $\mathscr{C}$ as the generator polynomial of $\mathscr{C}$. We use $d_{H}(\mathscr{C})$ or $d(\left[g(x)\right])$ to denote the minimum Hamming distance of $\mathscr{C}$. The parity check polynomial of $\mathscr{C}$ is $h(x)=(x^{n}-1)/g(x)$, whose reciprocal polynomial is $\tilde{h}(x)=x^{\deg(h(x))}h(x^{-1})$. The Euclidean dual of $\mathscr{C}$ is also a cyclic code, which has generator polynomial $g^{\perp_{e}}(x)=h_{0}^{-1}\tilde{h}(x)$, where $h_{0}$ is the constant term coefficient of $h(x)$. If $g^{\perp_{e}}(x)\mid g(x)$, then $\mathscr{C}$ is a Euclidean self-orthogonal code.
@@ II-C Quantum error-correction codes
Note: Throughout the paper, we agree on the following notation. All the indices $\ell$ of QC codes in the subsequent texts will be even and set $m=\ell/2$. For $f(x)=f_{0}+f_{1}x+\cdots+f_{n-1}x^{n-1}\in\mathbb{R}_{q,n}$, denote $\bar{f}(x)=x^{n}f(x^{-1})$. Let $\mathbf{0}$ and $\mathbf{0}_{n}$ denote a appropriate zero matrix and a zero vector of length $n$, respectively. Additionally, we denote $\Omega_{mn}=\left(\begin{array}[]{cc}\mathbf{0}&I_{mn}\\ -I_{mn}&\mathbf{0}\end{array}\right)$, where $I_{mn}$ is the identity matrix of order $mn$.
@@ III The necessary and sufficient conditions for QC codes with index even to be symplectic self-orthogonal
Let $g(x)$ be a polynomial in $\mathbb{R}_{q,n}$ that divides ${{x}^{n}}-1$, and let $f_{j}(x)$ be $\ell$ different polynomials in $\mathbb{R}_{q,n}$ for $10\leq j\leq\ell-$ 1, such that the greatest common divisor of $f_{0}(x),f_{1}(x),\ldots,f_{\ell-1}(x)$ and $\frac{x^{n}-1}{g(x)}$ is equal to $1$. If $\mathscr{C}$ is a QC code generated by $(\left[g(x){f}_{0}(x)\right]$, $\left[g(x){f}_{1}(x)\right],\ldots,\left[g(x){f}_{\ell-1}(x)\right])\in\mathbb{R}_{q,n}^{\ell}$, then $\mathscr{C}$ is called a $1$-generator QC code with index $\ell$. The generator matrix $G$ of $\mathscr{C}$ has the following form:
@@
Let $1\leq i\leq h$. Let $g_{i}(x)$ be a polynomial in $\mathbb{R}_{q,n}$ that divides ${{x}^{n}}-1$, and let $f_{i,j}(x)$ be $\ell$ different polynomials in $\mathbb{R}_{q,n}$ for $10\leq j\leq\ell-$ 1, such that the greatest common divisor of $f_{i,0}(x),$ $f_{i,1}(x),\ldots,f_{i,{\ell-1}}(x)$ and $\frac{x^{n}-1}{g_{i}(x)}$ is equal to $1$. If $\mathscr{C}$ is a QC code with $h$ generators given by: $(\left[g_{1}(x)f_{1,0}(x)\right],\left[g_{1}(x)f_{1,1}(x)\right],\ldots,\left[g_{1}(x)f_{1,\ell-1}(x)\right])$, $(\left[g_{2}(x)f_{2,0}(x)\right],$ $\left[g_{2}(x)f_{2,1}(x)\right],\ldots,\left[g_{2}(x)f_{2,\ell-1}(x)\right]),\ldots$, $(\left[g_{h}(x)f_{h,0}(x)\right],$ $\left[g_{h}(x)f_{h,1}(x)\right],\ldots,\left[g_{h}(x)f_{h,\ell-1}(x)\right])\in\mathbb{R}_{q,n}^{\ell}$, then $\mathscr{C}$ is an $h$-generator QC code with index $\ell$, whose generator matrix is
@@
If $\mathbf{c}^{\prime}=\mathbf{0}_{n}$ or $\mathbf{c}^{\prime\prime}=\mathbf{0}_{n}$, then we have $a(x)g(x)f_{0}(x)\equiv 0\pmod{x^{n}-1}$ or $a(x)g(x)f_{1}(x)\equiv 0\pmod{x^{n}-1}$. By Lemma 5, the nonzero half part of $\mathbf{c}$ is a codeword of $\mathscr{C}_{3}$ or $\mathscr{C}_{4}$. Therefore, it follows that $d_{s}(\mathscr{C})\leq\min\{d_{H}(\mathscr{C}_{3}),d_{H}(\mathscr{C}_{4})\}$.
@@
Secondly, if there is $\alpha$ in $\mathbb{F}_{q}^{*}$ such that $\mathcal{I}^{(\alpha)}(\mathscr{C})=\gcd(f_{0}(x)+\alpha f_{1}(x),h(x))\neq 1$, i.e., $\mathcal{S}\neq\emptyset$, then there exist $a(x)$ of degree less than $n-\deg(g(x))$ such that $a(x)(f_{0}(x)+\alpha f_{1}(x))g(x)\equiv 0\pmod{x^{n}-1}$, i.e., $\mathrm{Supp}(\mathbf{c}^{\prime})=\mathrm{Supp}(\mathbf{c}^{\prime\prime})$. For $q=2$, if $\mathrm{Supp}(\mathbf{c}^{\prime})=\mathrm{Supp}(\mathbf{c}^{\prime\prime})$, then we have $\mathbf{c}^{\prime}=\mathbf{c}^{\prime\prime}$ $\in\mathscr{C}_{1}\cap\mathscr{C}_{2}$. Therefore, $d_{s}(\mathscr{C})\geq d_{H}(\left[g(x)\operatorname{lcm}(f_{0}(x),f_{1}(x))\right])$. For $q>2$, we can only determine $d_{s}(\mathscr{C})\geq\max\{d_{H}(\mathscr{C}_{1}),d_{H}(\mathscr{C}_{2})\}$.
@@
With the above notation, let $\mathscr{C}$ be a $1$-generator QC code with generator $(\left[f_{0}(x)g(x)\right],\left[f_{1}(x)g(x)\right])$. If $\gcd(f_{0}(x),g(x))=1$, then the symplectic dual code $\mathscr{C}^{\perp_{s}}$ of $\mathscr{C}$ is a 2-generator QC code, whose generator matrix is
@@
where $(F_{0},F_{1})$ and $(\mathbf{0},G_{g^{\perp_{e}}})$ are the generator matrices of $1$-generator QC codes determined by $(\left[\bar{f}_{0}(x)\right],\left[\bar{f}_{1}(x)\right])$ and $(\left[0\right],\left[g^{\perp_{e}}(x)\right])$, respectively.
@@
Let $\mathscr{C}^{\circ}$ denote the code generated by $G^{\perp_{s}}$. Then any codeword of $\mathscr{C}$ and $\mathscr{C}^{\circ}$ has the form $\mathbf{c_{1}}=(\left[a(x)f_{0}(x)g(x)\right],$ $\left[a(x)f_{1}(x)g(x)\right])$ and $\mathbf{c_{2}}=(\left[b(x)\bar{f}_{0}(x)\right],\left[b(x)\bar{f}_{1}(x)+c(x)g^{\perp_{e}}(x)\right])$, where $a(x),b(x),c(x)\in\mathbb{R}_{q,n}$. Using a similar approach of Theorem 2, it is easy to prove that $\left\langle\mathbf{c_{1}},\mathbf{c_{2}}\right\rangle_{s}=\mathbf{c_{1}}\cdot\Omega_{mn}\cdot\mathbf{c_{2}}^{T}=0$. Therefore, $\mathscr{C}$ and $\mathscr{C}^{\circ}$ are mutually symplectic orthogonal.
@@
According to Definition 1, $\gcd(\bar{f}_{0}(x),\bar{f}_{1}(x),g^{\perp_{e}}(x))=1$. Given that $\gcd(f_{0}(x),g(x))=1$, we have $\gcd(\bar{f}_{0}(x),\tilde{g}(x))=1$. As $\tilde{g}(x)g^{\perp_{e}}(x)=x^{n}-1$, $\gcd(\bar{f}_{0}(x),\bar{f}_{1}(x),x^{n}-1)=1$ and $\operatorname{lcm}\left(g^{\perp_{e}}(x),\frac{x^{n}-1}{\gcd(x^{n}-1,\bar{f}_{0}(x))}\right)\equiv 0\pmod{x^{n}-1}$ hold. This implies that, the rank of $(F_{0},F_{1})$ is $n$ and the rows of $B$ and $G_{g^{\perp_{e}}}$ are linearly independent. Therefore, the rank of $G_{1}^{\perp_{s}}$ is $n+\deg(g(x))$. Thus, $\mathscr{C}$ and $\mathscr{C}^{\circ}$ are symplectic dual codes. ∎
@@
With the above notation, let $\mathscr{C}$ be a $1$-generator QC code with generator $(\left[f_{0}(x)g(x)\right],\left[f_{1}(x)g(x)\right])$ satisfying $\gcd(f_{0}(x),g(x))=1$. For $\alpha\in\mathbb{F}_{q}^{*}$, define $\mathcal{I}^{(\alpha)}_{\perp}(\mathscr{C})=\gcd(\bar{f}_{1}(x)+\alpha\bar{f}_{0}(x),g^{\perp_{e}}(x))$. Let $\mathcal{S}_{\perp}=\{\mathcal{I}^{(\alpha)}_{\perp}(\mathscr{C})\mid\mathcal{I}^{(\alpha)}_{\perp}(\mathscr{C})\neq 1,\forall\alpha\in\mathbb{F}_{q}^{*}\}$.
Suppose that $\mathscr{C}_{0}$, $\mathscr{C}_{1}$, $\mathscr{C}_{2}$, $\mathscr{C}_{3}$, $\mathscr{C}_{4}$, $\mathscr{C}_{5}$, $\mathscr{C}_{6}$ and $\mathscr{C}_{7}$ are cyclic codes determined by $\left[g^{\perp_{e}}(x)\right]$, $\left[\bar{f}_{0}(x)\right]$, $\left[\gcd(\bar{f}_{1}(x),g^{\perp_{e}}(x))\right]$, $\left[\frac{x^{n}-1}{\gcd(x^{n}-1,\bar{f}_{1}(x))}\right]$, $\left[\frac{x^{n}-1}{\gcd(x^{n}-1,\bar{f}_{0}(x))}\right]$, $\left[\gcd\left(\frac{x^{n}-1}{\gcd(x^{n}-1,\bar{f}_{0}(x))},g^{\perp_{e}}(x)\right)\right]$, $\left[\operatorname{lcm}(\bar{f}_{0}(x),\gcd(\bar{f}_{1}(x),g^{\perp_{e}}(x))\right]$ and $\left[\bar{f}_{0}(x)\frac{g^{\perp_{e}}(x)}{\gcd(\bar{f}_{1}(x),g^{\perp_{e}}(x))}\right]$, respectively.
@@
Let $t=(n-\deg(\gcd(x^{n}-1,\bar{f}_{1}(x)))-\deg(\gcd(x^{n}-1,\bar{f}_{0}(x))))$ and $\mathscr{C}_{8}$ be the cyclic code generated by $\left[\bar{f}_{1}(x)\right]$. Since $\gcd(f_{0}(x),g(x))=1$, by Theorem 5, the symplectic dual of $\mathscr{C}$ is a 2-generator QC code with generator matrix
@@
If $a(x)\bar{f}_{0}(x)\not\equiv 0\pmod{x^{n}-1}$ and $a(x)\bar{f}_{1}(x)\not\equiv 0\pmod{x^{n}-1}$, there are the following scenarios to discuss.
@@
Thirdly, if there exists $\alpha$ in $\mathbb{F}_{q}^{*}$ such that $\gcd(\bar{f}_{1}(x)+\alpha\bar{f}_{0}(x),g^{\perp_{e}}(x))\neq 1$, i.e., $\mathcal{S}_{\perp}\neq\emptyset$, then there may exist $\mathrm{Supp}(\left[a(x)\bar{f}_{0}(x)\right])=\mathrm{Supp}(\left[a(x)\bar{f}_{1}(x)+b(x)g^{\perp_{e}}(x)\right])$. Thus, for $q>2$, we can only deduce $d_{s}(\mathscr{C}^{\perp_{s}})\geq\max\left\{d_{H}(\mathscr{C}_{1}),d_{H}(\mathscr{C}_{2})\right\}$. But for $q=2$, we have $d_{s}(\mathscr{C}^{\perp_{s}})\geq d_{H}(\left[\operatorname{lcm}(\bar{f}_{0}(x),\gcd(\bar{f}_{1}(x),g^{\perp_{e}}(x)))\right])$.
For the case of $\mathrm{Supp}(\left[a(x)\bar{f}_{0}(x)\right])\neq$ $\mathrm{Supp}(\left[a(x)\bar{f}_{1}(x)+\right.$ $\left.b(x)g^{\perp_{e}}(x)\right])$, we have Equation (25) holds.
@@
Remark 3
The new binary QECCs presented in this paper have been updated in Grassl’s code table [12]. Additionally, after we completed this paper, Grassl et al. [35] systematically constructed a large number of good QECCs using QC codes, and extended the quantum code table to $q\leq 8$ and $n\leq 100$.